LASCO Activity Report for
October 31, 2000 SOHO Science Daily Meeting

Planner: S. Plunkett, K. Schenk

Event times are first frame seen in C2 camera unless otherwise noted.

Points Of Interest

None Today.....


	Lasco calibrations continue.....
	Synoptic program resumed at 12:30 UT.

18:06 UT  S East	Large bright loop front event.  Interior core
			with loops.  Continued outflow.
                20:06:05 UT C2 Running Difference Image. 

18:34 UT  N West	Ragged front CME, fast, small, disipates quickly.

21:26 UT  N West	Faint blob front.

22:06 UT  East Limb	Narrow jet.


03:50 UT SS West	Rising multiple loops over limb, very slow.
			Still visible at 07:50 before DSN data gap.
                07:50:05 UT C2 Running Difference Image. 

Web curator: K.M.Schenk
NASA Goddard Space Flight Center
Solar Physics Branch / Code 682
Greenbelt, MD 20771